OpenStudy (anonymous):

Choose the correct simplification of the expression (2x2y6z5)(5x4y5z3).

OpenStudy (anonymous):

10 x8y30z15 7 x6y11z8 7x8y30z15 10x6y11z8 I believe its a

OpenStudy (anonymous):

@phi how long did it take you to achieve the smart score and become a moderater

OpenStudy (anonymous):

@iGreen

OpenStudy (phi):

when you multiply you can switch the order for example \[ (2x^2y^6z^5)(5x^4y^5z^3) = 2\cdot 5 \cdot x^2x^4y^6y^5z^5z^3 \] hopefully you know 2*5 is 10 so the answer has a 10 in it. you should know that x^2 is short for x*x (x times itself 2 times) and x^4 is x*x*x*x (times itself 4 times) so x^2 * x^4 is x*x * x*x*x*x (x times itself 6 times) of course, it's easier to write x^6 than x*x * x*x*x*x same idea for the y's and z's

OpenStudy (phi):

you got the right answer